OpenStudy (anonymous):

Simplify: 3 √27(x+1)^3

OpenStudy (anonymous):

\[3*\sqrt(27*(x+1)^3)=3*\sqrt([3*3]*3*[(x+1)*(x+1)]*(x+1))\] \[3*3*[(x+1)*\sqrt(3(x+1))=(9x+9)\sqrt(3x+3)\]

OpenStudy (anonymous):

thank you. I find this simplifying insanely difficult. any tips?

OpenStudy (anonymous):

Try to break down everything that you can. Anything with an exponent (x+a)^4 can be broken down my multiplying it multiple times. Any even numbers can be factored by 2. Just try to break everything down as much as possible that way you can visualize what you can factor easier.

OpenStudy (anonymous):

ok great..what does the 3 on the far left mean?